OpenStudy (oliviagggg):

MEDAL AND FAN PLEASE HELP A 12-foot flagpole casts a nine-foot shadow. What is the measure from the top of the shadow to the top of the flag pole? 15 21 3 6

OpenStudy (anonymous):

one sec

OpenStudy (faiqraees):



OpenStudy (faiqraees):

Using pythagoras theorm a^2 = b^2+c^2 x^2 = 12^2+9^2 Can you solve for x?

OpenStudy (mathstudent55):



OpenStudy (oliviagggg):

i got 15

OpenStudy (anonymous):

I got 15 as well

OpenStudy (oliviagggg):

Thanks :)

OpenStudy (anonymous):

do one of you mind helping me with a problem?

OpenStudy (oliviagggg):

put their names in your question

OpenStudy (anonymous):

okay thank you